OpenStudy (anonymous):

8/9 - 12/9 = 7/8 - (-3/16) = -1/12 - 3/4 =

OpenStudy (anonymous):

8/9-12/9=8-12/9 Answer !!!

OpenStudy (anonymous):

-4/9 4/9 -4/18 those are the answer chose's doe

OpenStudy (anonymous):

-4/9

OpenStudy (anonymous):

what about the others?

OpenStudy (anonymous):

heh !

OpenStudy (anonymous):

7/8 - (-3/16) = 7/8+3/16=14+3/16=17/16 -1/12 - 3/4 =-1-9/12=-10/12